### v2

# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

Abstract.
We prove that the nonnegative $3$-intermediate Ricci curvature and uniformly positive $k$-triRic curvature implies rigidity of complete noncompact two-sided weakly stable minimal hypersurfaces in a Riemannian manifold $(X^{5},g)$ with bounded geometry. The nonnegativity of $3$-intermediate Ricci curvature can be replaced by nonnegative Ricci and biRic curvature. In particular, there is no complete noncompact stable minimal hypersurface in a closed $5$-dimensional manifold with positive sectional curvature. It extends result of Chodosh-Li-Stryker [to appear in J. Eur. Math. Soc (2024)] to $5$-dimensions. We also prove some nonexistence results of complete noncompact constant mean curvature hypersurfaces with finite index in hyperbolic space $\mathbb{H}^{5}$.
Key words and phrases:
1. Introduction
An immersed complete two-sided stable minimal hypersurface $M^{n}$ in $(X^{n+1},g)$ has zero mean curvature and satisfies the inequality
(1.1) $\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for any $\psi\in C^{\infty}_{c}(M)$, where $A_{M}$ is the second fundamental form of the immersion, $\eta$ is a unit normal vector on $M$ and $\operatorname{Ric}_{X}$ is the ambient Ricci curvature. In a similar manner of stable geodesic in $2$-dimensions, stable minimal hypersurfaces is a powerful tool to study the ambient geometry of $(X^{n+1},g)$ in higher dimensions. Assuming $M$ is compact, basic arguments show that
  1. (i)
    When $\operatorname{Ric}_{X}>0$, there are no closed two-sided stable minimal hypersurfaces. If $\operatorname{Ric}_{X}\geq 0$, then $M$ is totally geodesic and $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$; see [Simons-minimal-varieties].
  2. (ii)
    If the ambient scalar curvature is strictly positive, then $M$ also admits a metric of positve scalar curvature [Schoen-Yau-incompressible-minimal-surfaces, Schoen-Yau-PSC]. In three dimension, each component of $M$ must be a $2$-sphere or projective plane.
Now, if one assumes that $M^{n}$ is non-compact with respect to the induced metric, it becomes very complicated and subtle. The theory has been well-developed in $3$-dimensions: If $M^{2}\hookrightarrow(X^{3},g)$ is a complete noncompact two-sided stable minimal immersion, it holds that
  1. (i)
    When $R_{X}\geq 0$, $(M,g)$ is conformal to either a plane or a cylinder [Fischer-Colbrie-Schoen-The-structure-of-complete-stable]. In the latter case, $M$ is totally geodesic, intrinsically flat and $R_{X}\equiv 0$ along $M$.
  2. (ii)
    When $\operatorname{Ric}_{X}\geq 0$, $(M,g)$ is totally geodesic, intrinsically flat and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$ [Schoen-Yau-positive-Ricci-curvature]. Moreover, either $X^{3}$ is diffeomorphic to $\mathbb{R}^{3}$ or its universal cover is isometric to the Riemann product $M^{2}\times\mathbb{R}$. [Liu-nonnegative-Ricci-curvature].
  3. (iii)
    When $R_{X}\geq 1$, $M$ must be compact [Schoen-Yau-black-hole].
Recently, Chodosh, Li and Stryler developed such theory in $4$-dimensions by combining nonnegativity of $2$-intermediate Ricci curvature with strict positivity of scalar curvature.
Theorem 1.1 ([chodosh-li-stryker]).
If $(X^{4},g)$ is a complete $4$-dimensional Riemannian manifold with weakly bounded geometry, satisfying
$\operatorname{Ric}^{X}_{2}\geq 0\ \text{and}\ R_{X}\geq\delta>0,$
where $\operatorname{Ric}_{2}^{X}$ is the $2$-intermediate Ricci curvature (Definition 1.7). Then any complete two-sided stable minimal immersion $M^{3}\rightarrow(X^{4},g)$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M$.
It is natural to ask what happens in $5$-dimensions. For example, a basic question is to determine which ambient curvature conditions imply compactness or non-existence of complete stable minimal hypersurfaces. As discussed above in $3$-manifolds, $R_{X}\geq 1$ implies compactness and $\operatorname{Ric}_{X}>0$ implies non-existence. Two examples constructed in [chodosh-li-stryker][Example 1.1,1.2] can be easily adapted to show that neither of these results can hold in $5$-dimensions.
Note that for most of aforementioned results, it is essential that no assumption is made on properness or volume growth of $M$. The result in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because $\operatorname{Ric}^{X}_{2}\geq 0$ and $R_{X}>\delta$ are tailor to obtain one-endness and volume growth of hypersurfaces in $4$-dimensions. Hence, exploring alternative curvature conditions to maintain one-endness and volume growth in higher dimensions is one of the main topics in this paper.
1.1. Main results
The purpose of this article is twofold. First, we extend results in [chodosh-li-stryker] to $5$-dimensions under different curvature assumptions and weak stability of minimal hypersurface. Second, we prove that there are no complete noncompact constant mean curvature (CMC) hypersurfaces with (averaged) mean curvature $H_{M}>\frac{\sqrt{65}}{8}$, finite index and universal subexponential end growth in hyperbolic space $\mathbb{H}^{5}$, as a special case of a more general nonexistence result.
Before discussing our main results, we recall some definitions and introduce some notions.
Definition 1.2.
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. An immersed CMC hypersurface (two-sided if it is minimal) $M^{n}$ in $X$ is weakly stable if
$\int_{M}\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)\psi^{2}\leq\int_{M}|\nabla\psi|^{2}$
for $\psi\in C_{c}^{\infty}(M)$ satisfying $\int_{M}\psi=0.$ We call it stable if we do not assume $\int_{M}\psi=0$. Moreover, $M$ is said to have finite index if the operator $\Delta_{M}+|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)$ has finite number of negative Dirichlet eigenvalues in any compact subset of $M$.
Definition 1.3.
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. Given a local orthonormal basis $\{e_{i}\}_{i=1}^{n+1}$ on $T_{p}X$ and $k>0$, we define $k$-weighted triRicci curvature as
(1.2) $k\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})=k\operatorname{Ric}_{X}(e_{i},e_{i})+\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}),$
where $\operatorname{biRic}^{e_{i}^{\perp}}(e_{j},e_{l})$ is given by
$\operatorname{biRic}_{e_{i}^{\perp}}(e_{j},e_{l}):=\sum_{m\neq i}\operatorname{Rm}^{X}_{mjjm}+\sum_{m\neq i}\operatorname{Rm}^{X}_{mllm}-\operatorname{Rm}^{X}_{jllj},\quad j\neq l$
i.e., we take the partial contraction on the subspace $e_{i}^{\perp}:=\operatorname{span}\{e_{l}\}_{l\neq i}$ normal to $e_{i}$.
Remark 1.4.
When $n=2$ or $3$, $\operatorname{1-triRic}$ is just the half of the scalar curvature of $X$. In particular, $1$-triRic curvature is the $3$-intermediate curvature $\mathcal{C}_{3}$ defined by Brendle-Hirsch-Johne[brendlegeroch'sconjecture].
Remark 1.5.
Direct calculation yields
$\begin{split}&\sum_{i=1}^{n+1}\sum_{j\neq i,l\neq i}k\operatorname{-triRic}_{X}(e_{i};e_{j},e_{l})\\ &=\sum_{i=1}^{n+1}\left(kn(n-1)\operatorname{Ric}_{X}(e_{i},e_{i})+(2n-3)\sum_{j\neq i,l\neq i}\operatorname{Rm}^{X}_{jllj}\right)\\ &=kn(n-1)\sum_{i=1}^{n+1}\operatorname{Ric}_{X}(e_{i},e_{i})+(2n-3)(n-1)\sum_{j,l=1}^{n+1}\operatorname{Rm}^{X}_{jllj}\\ &=((k+2)n-3)(n-1)R_{X}.\end{split}$
In particular, positive $k$-weighted triRicci curvature implies that $X$ has positive scalar curvature. On the converse, $\Sigma^{2}\times S^{3}(5/12)$ has positive scalar curvature but not always positive $1$-weighted triRicci curvature where $\Sigma^{2}$ is a closed hyperbolic surface and $S^{3}(5/12)$ is a sphere with constant curvature $5/12$.
Remark 1.6.
When $(X^{n+1},g)$ is the hyperbolic space, $k\operatorname{-triRic}_{X}=3-(k+2)n.$
Definition 1.7.
Let $(X^{n+1},g)$, $n\geq 2$, be a Riemannian manifold. The $(n-1)$-intermediate Ricci curvature of $X$ in the direction $\mu$, is defined as
$\operatorname{Ric}^{X}_{n-1}(e_{i};\mu)=\sum_{j\in\{1,\cdots,n\}\setminus i}\operatorname{Rm}_{X}(e_{j},\mu,\mu,e_{j})$
where $\{e_{1},\cdots,e_{n},\mu\}$ is any orthonormal basis of $T_{p}X$.
Throughout this article, for any tensor $T$ acting on tangent vectors we say $T\geq C$ for a constant $C$ if $T(e_{i},\cdots)\geq C$ for any unit vectors $e_{i},\cdots\in T_{p}X$ and any $p\in X.$ Note that $\operatorname{Ric}_{n-1}^{X}\geq 0$ implies that $1\operatorname{-triRic}_{X}\geq 0$.
We now turn to our first main result.
Theorem 1.8.
If $(X^{5},g)$ is a complete $5$-dimensional Riemannian manifold with bounded geometry, satisfying
$\operatorname{Ric}^{X}_{3}\geq 0\ \text{and}\ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2].$
Then any complete two-sided weakly stable minimal immersion $M^{4}\rightarrow X^{5}$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M.$
It is possible to weaken non-negativity of $\operatorname{Ric}_{3}^{X}$ to a mixed curvature condition. Note that $\operatorname{Ric}_{3}^{X}\geq 0$ implies $\operatorname{Ric}_{X}\geq 0$ and $\operatorname{biRic}_{X}\geq 0$, where $\operatorname{biRic}_{X}$ is the bi-Ricci curvature (see Definition 2.2 in detail). In the Appendix, we provide a (one parameter family) closed $5$-dimensional Riemannian manifold $(X^{5},g)$, satisfying
$\operatorname{Ric}_{X}>0,\ \,\operatorname{biRic}_{X}>0,\ \,\text{and}\ \ 2\operatorname{-triRic}_{X}\geq\delta>0,$
while $\operatorname{Ric}_{3}<0$ somewhere. As a matter of fact, the result in Theorem 1.8 holds with $\operatorname{Ric}_{3}^{X}\geq 0$ replaced by $\operatorname{Ric}_{X}\geq 0$ plus $\operatorname{biRic}_{X}\geq 0$.
Theorem 1.9.
If $(X^{5},g)$ is a complete $5$-dimensional Riemannian manifold with bounded geometry, satisfying
$\operatorname{Ric}_{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2].$
Then any complete two-sided weakly stable minimal immersion $M^{4}\rightarrow X^{5}$ is totally geodesic and has $\operatorname{Ric}_{X}(\eta,\eta)\equiv 0$ along $M.$
Let us make some comments here. The first one is that if $(X^{5},g)$ is a manifold with bounded geometry satisfying $\operatorname{Ric}_{X}>0$, $\operatorname{biRic}_{X}\geq 0$ and $k\operatorname{-triRic}_{X}\geq\delta>0\ \text{for}\ k\in[3/2,2]$, then there is no complete two-sided weakly stable minimal immersion. This follows directly from Theorem 1.9. The second one is that there is no complete two-sided weakly stable minimal immersion in manifold $X$ with $\operatorname{Ric}_{X}>0$ and $\operatorname{biRic}_{X}\geq\delta>0$. Indeed, uniformly positive $\operatorname{biRic}_{X}$ implies that a complete two-sided weakly stable minimal immersion must be compact [shenyingyerugang]. Then the standard argument leads to contradiction. The third one is that Theorem 1.9 also implies that there is no two-sided weakly stable minimal immersion in $(\mathbb{S}^{5},g_{can})$ (previously proved by [catino]). Lastly, the proof of above theorems also work for weakly stable CMC immersion.
The second result is about the nonexistence of complete noncompact CMC hypersurfaces with finite index. We first recall universal subexponential end growth condition for a noncompact manifold $M$ (possibly with noncompact boundary).
Definition 1.10 ([Hong24]).
For a fixed $R>0$ and $p\in M$, define $N_{M,k}$ to be the number of components of $M\cap(B_{(k+1)L+R}\setminus B_{kL+R})$:
$N_{M,k}:=\#\{M\cap(B_{(k+1)L+R}\setminus B_{kL+R})\}.$
We say that $M$ has subexponential end growth if
$\limsup_{k\rightarrow\infty}\frac{N_{M,k}}{e^{Ck}}=0$
for any constant $C>0.$
Note that if $M$ has boundary, $M\cap(B_{(k+1)L+R}\setminus B_{kL+R})$ contains two types of components, depending on whether they have boundary portion belonging to $\partial M.$
Definition 1.11 ([Hong24]).
A complete noncompact manifold $M$ has universal subexponential end growth if there is one connected component of $M\setminus B_{R_{0}}$ whose universal covering satisfies subexponential end growth for some fixed $R_{0}>0$.
Theorem 1.12.
There is no complete noncompact constant mean curvature hypersurface $M$ in hyperbolic space $\mathbb{H}^{5}$ satisfying universal subexponential end growth condition,
$H_{M}>\frac{\sqrt{65}}{8}\ \ \text{and}\ \ \operatorname{Ind}<\infty.$
The question of whether a complete constant mean curvature hypersurface in $\mathbb{H}^{n+1}$, where $n\geq 4$, with finite index and $H>1$ must be compact remains open in general. Previously, Da Silveira confirmed this for the case $n+1=3$ in [daSilveira-Stability-of-complete], while the positive answer for $n+1=4$ was provided by the first author in [Hong24] under the universal subexponential end growth condition. Partial progress has been made under various stronger or additional assumptions. For instance, in dimension $n+1=5$ [chengxufiniteindex, dengqintao], an assumption was imposed on mean curvature: $H>\frac{\sqrt{175}}{\sqrt{148}}$. Thus, our theorem presents a contribution when $H\in(\frac{\sqrt{65}}{8},\frac{\sqrt{175}}{\sqrt{148}}]$. We remark that the universal subexponential end growth condition in above theorem can be replaced by finite first Betti number and finite ends. Let us mention that when zero volume entropy is additionally assumed, a complete positive solution was achieved in all dimensions [zeroentropy].
Indeed, we can prove a nonexistence result under more general settings; see Theorem 5.2 in detail. Consequently, Theorem 1.12 follows from Theorem 5.2 directly.
1.2. Main ideas
The strategy in this article follows from that in [chodosh-li-stryker] and [Hong24]. A common feature in the proofs of Theorem 1.8 and 1.12 is the estimate on volume growth of ends, using $\mu$-bubbles, a technique introduced by Gromov [gromovmububble]. The method in [chodosh-li-stryker] cannot extend to $5$-dimensions directly because one cannot control curvature terms from the second fundamental form, the ambient Ricci curvature and the positive Jacobi function in the second variation of the warped $\mu$-bubble constructed there. The novelty of our method is to introduce a parameter (the exponent of the warped function) which brings more freedom to exploit curvature conditions. The same strategy has been applied successfully in the study of CMC hypersurfaces in $\mathbb{H}^{4}$ [Hong24].
Another interesting feature of Theorem 1.9 and Theorem 5.2 is the combination of different curvature conditions which enter the proof at different places. For completeness, we present the overall strategy as follows.
In Theorem 1.9, consider $M^{4}\hookrightarrow(X^{5},g)$ a complete two-sided weakly stable minimal hypersurface.
  1. (i)
    To show the vanishing of $A_{M}$ and $\operatorname{Ric}_{X}(\eta,\eta)$, on the contrary we show that if $A_{M}+\operatorname{Ric}_{X}(\eta,\eta)$ is strictly positive somewhere, we can find a sequence of compactly supported test functions $\varphi_{i}$ such that $\int_{M}|\nabla\varphi_{i}|^{2}=o(1)$, thus contradicting the weak stability.
  2. (ii)
    In the same spirit in [Cao-Shen-Zhu-infinitevolume], the fact that $M$ has at most one nonparabolic end follows from the Schoen–Yau inequality (just uses $\operatorname{biRic}_{X}\geq 0$) and the infinite volume of $M$ (guaranteed by bounded geometry). If we assume $\operatorname{Ric}^{X}_{3}\geq 0$, bounded geometry condition can be weaken to weakly bounded geometry condition; see [chodosh-li-stryker][Theorem 4.1] in detail.
  3. (iii)
    The construction of a good test function along the nonparabolic end is the main issue in the remaining proof. This is where the positivity of $k\operatorname{-triRic}$ comes in. By conformal change, it is well known [Schoen-Yau-PSC] that a stable minimal hypersurface admits a metric with positive scalar curvature if the ambient metric does. It is interesting to understand if the positivity of $k\operatorname{-triRic}$ can be inherited by hypersurfaces in some sense.
    As mentioned in [chodosh-li-stryker], it is difficult to prove the linear volume growth of the nonparabolic end. Similarly, we consider an exhaustion $\{\Omega_{l}\}_{l=1}^{\infty}$ of $E$ and use the theory of $\mu$-bubbles to show $\partial\Omega_{l}\cap E$ has controlled diameter. To control curvature terms in the second variation of the $\mu$-bubble, we introduce a parameter $k$ as the exponent of the warped function. Suitable choice of $k$ yields the desired diameter estimate; see Lemma 4.6 in detail. This is the step that needs the assumption of strictly positive $k\operatorname{-triRic}$.
  4. (iv)
    We need to upgrade diameter control to volume control which needs a lower bound on the intrinsic Ricci curvature of $M$. It is achieved by combining $\operatorname{biRic}_{X}\geq 0$, curvature estiamtes in Lemma 2.4 and bounded geometry.
  5. (v)
    Finally, we need to construct balanced test function for weak stability in step (i). This is achievable due to the facts that we have at most one nonparabolic end and the nonparabolic end satisfies almost linear volume growth. In fact, we can prove the following general result which is of independent interest.
Proposition 1.13.
Let $M$ be a complete manifold with infinite volume and finite nonparabolic ends $\{E^{k}\}_{k=1}^{k_{0}}$. For a fixed scale $L$ and each $k$, let $E^{k}_{l}$ be the nonparabolic component of $E^{k}\setminus B_{(l+1)L}(p)$. Suppose that each $E^{k}$ satisfies almost linear volume growth, i.e., $E_{l}\cap B_{(l+1)L}(p)\leq C$ for a universal constant $C$. Then there exists a compact supported Lipschitz function $\varphi$ such that
$\int_{M}\varphi=0\ \text{and}\ \ -\int_{M}\varphi L(\varphi)<0$
for any operator $L=\Delta_{M}+q(x)$ where $q(x):M\rightarrow\mathbb{R}_{\geq 0}$ is not identically zero.
The proof of Theorem 5.2 follows the same strategy in [Hong24] by the first author. We sketch it as follows: The finite index assumption implies that any end $E$ of $M$ is stable outside a compact set. Combining the condition 5.1 with the stability inequality yields the exponential volume growth of the universal covering $\widetilde{E}$, which contradicts with the almost linear growth implied by Lemma 4.7. This is where condition (5.2) and universal subexponential end growth condition enter the proof.
We end the introduction with the following conjecture.
Conjecture: There is no complete noncompact constant mean curvature hypersurface $M$ in hyperbolic space $\mathbb{H}^{5}$ satisfying
$\frac{\sqrt{175}}{\sqrt{148}}>H_{M}>1\ \ \text{and}\ \ \operatorname{Ind}<\infty.$
1.3. Organization
In Section 2, we recall some preliminaries on curvature estimates and the ends of non-compact manifolds. In Section 3, we prove the one-end theorem under $\operatorname{biRic}_{X}\geq 0$ and bounded geometry assumption. We discuss the balanced warped $\mu$-bubble exhaustion result and the volume growth estimates in Section 4, and prove main results in Section 5. Appendix contains an example relevant to Theorem 1.9.
1.4. Acknowledgements
The first author is supported by a Science grant (YA24XKRC00010) by Beijing Jiaotong University. We would like to thank Prof. Xianzhe Dai, Prof. Martin Li, Prof. Guofang Wei and Prof. Rugang Ye for their interest in this work.
2. Preliminary
2.1. Curvature conditions
If $(X^{n+1},g)$ is a Riemannian manifold, we denote the Riemann curvature tensor, the sectional curvature, the Ricci curvature and the scalar curvature by $\operatorname{Rm}_{X}$, $\sec_{X}$, $\operatorname{Ric}_{X}$ and $R_{X}$ respectively. Let $M\hookrightarrow X^{n+1}$ be an immersion. We write $\operatorname{Ric}_{M}$ and $R_{M}$ to denote the Ricci curvature and scalar curvature of the pullback metric on $M$. We set the mean curvature to be the trace of the second fundamental form of the immersion.
Definition 2.1.
A manifold $X^{n+1}$ has bounded geometry if there exist positive real numbers $\delta$ and $\lambda$ such that on $X^{n+1}$
$\sec_{X}\leqslant\delta^{2},\quad\operatorname{inj}_{X}\geqslant\lambda$
where $\operatorname{inj}_{X}$ is the injectivity radius.
Definition 2.2.
Let $(X^{n+1},g)$ be a Riemannian manifold and $\{e_{i}\}_{i=1}^{n+1}$ be an orthonormal basis for $T_{p}X$, $p\in X$. The bi-Ricci curvature is defined by
$\operatorname{biRic}_{X}(e_{1},e_{2}):=\sum_{i=2}^{n+1}\operatorname{Rm}_{X}(e_{1},e_{i},e_{i},e_{1})+\sum_{j=3}^{n+1}\operatorname{Rm}_{X}(e_{2},e_{j},e_{j},e_{2}).$
This concept was used by Shen-Ye [shenyingyerugang] to study diameter of stable minimal submanifolds.
Lemma 2.3.
Let $(X^{5},g)$ be a $5$-dimensional Riemannian manifold with bounded geometry and $\operatorname{biRic}_{X}\geq 0$. If $M^{4}\hookrightarrow(X^{5},g)$ is an immersed CMC hypersurface, then
$\operatorname{Ric}_{M}\geqslant-Cg,$
where the positive constant $C$ depends on $\sec_{X}$, $H_{M}$ and $|A_{M}|$.
Proof.
Note that bounded geometry and nonnegative bi-Ricci condition imply that
$|\sec_{X}|\leqslant 6\delta^{2}.$
Choose an orthonormal basis $\{e_{i}\}_{i=1}^{5}$ of $X^{5}$ at a point $p\in M$ with $e_{5}$ normal to $M$. The traced Gauss equation gives
$\operatorname{Ric}_{M}(e_{1},e_{1})=\sum_{i=2}^{4}\operatorname{Rm}^{X}_{i11i}+A_{M}(e_{1},e_{1})H_{M}-\sum_{j=1}^{4}A_{M}(e_{1},e_{j})^{2},$
which yields the desired result. ∎
2.2. Finite index and curvature estimates
Recently, Chodosh, Li, Minter and Stryker proved that a two-sided complete stable minimal hypersurface $M^{4}\hookrightarrow\mathbb{R}^{5}$ is flat [Chodosh-Li-Minter-Stryker]. A standard blow-up argument yields curvature estimates for stable CMC hypersurfaces in non-compact $5$-manifolds.
Lemma 2.4.
Let $(X^{5},g)$ be a complete Riemannian manifold with bounded geometry and $\operatorname{biRic}\geqslant 0$. Then there is a constant $C$ such that every noncompact stable CMC hypersurface (two-sided if it is minimal) $M^{4}\hookrightarrow(X^{5},g)$ satisfies
$\sup_{q\in\Omega}|A_{M}(q)|\min\{1,d_{M}(q,\partial_{1}\Omega)\}\leqslant C$
for any compact subset $\Omega\subset M$ where $\partial_{1}\Omega=\partial\Omega\cap\mathring{M}.$
Proof.
Note that bounded geometry and $\operatorname{biRic}\geqslant 0$ implies the boundness of $\sec_{X}$. Proposition D.1 in [chodosh-li-stryker] shows that $Q$-weakly bounded geometry follows from $|\sec_{X}|\leqslant K$. Then, the proof is similar to that of [chodosh-li-stryker][Lemma 2.4] and [Hong24][Lemma 4.5] with minor modification, see also [Rosenberg-Souam-Toubiana]. As indicated in the proof of Lemma 4.5 in [Hong24], we only need to take care of the case when $M$ is not geodesically complete. ∎
2.3. Parabolicity and nonparabolicity
In this subsection, we recall some preliminary results [Hong24] on parabolicity and nonparabolicity of manifolds with noncompact boundary; see [chodosh-li-stryker] for detailed discussion on manifolds without boundary.
Let $M^{n}$ be a noncompact Riemannian manifold with boundary $\partial M$. Denote the outer unit normal vector of $\partial M$ by $\tau$. Choose a large compact subset $K$ containing $p\in\partial M$ such that the complement $M\backslash K$ has many disjoint components which are called ends. Boundary ends are those sharing boundary portion with $\partial M$ while interior ends only have boundary lying in $\partial K$. In dimension two, these two concepts are used to study the structure of capillary surfaces by the first author [hongcrelle].
Definition 2.5 ([Wu23]).
Denote a boundary end by $E$ and $\partial E$ consists of two parts: $\partial_{1}E:=\partial E\cap\partial K$ and $\partial_{0}E:=\partial E\cap\partial M$. $E$ is parabolic if there is no positive harmonic function $f\in C^{2,\alpha}(E)$, $10<\alpha<$ 1, so that
$f|_{\partial_{1}E}=1,\quad\partial_{\tau}f|_{\partial_{0}E}=0,\quad f|_{\mathring{E}}<1.$
Otherwise, we say that $E$ is nonparabolic and such $f$ is said to be a barrier function of $E$.
Remark 2.6.
The claimed regularity can be guaranteed by perturbing the intersection angle btween $\partial_{0}E$ and $\partial_{1}E$ to lie in $(0,\frac{\pi}{8})$ [Wu23]. In the sequel, we always assume that $\partial_{0}E$ meets $\partial_{1}E$ with a constant angle in $(0,\frac{\pi}{8})$.
We conclude this section by two useful characterizations of nonparabolicity.
Proposition 2.7 ([Wu23]Lemma 4.8).
A boundary end $E$ is nonparabolic if and only if the sequence of positive harmonic functions defined on $E_{i}:=E\cap B_{R_{i}}$, $R_{i}\to\infty$, satisfying
$f_{i}|_{\partial_{1}E}=1,\quad\partial_{\tau}f_{i}|_{\partial_{0}E}=0,\quad f_{i}|_{\partial B_{R_{i}}\cap E}=0$
converges uniformly and locally on compact subsets of $E$ to a barrier function $f$. Moreover, the barrier is minimal among all harmonic functions satisfying mixed boundary conditions, and has finite Dirichlet energy.
If the end $E$ is parabolic, the sequence constructed above converges to constant function $1$ and the Dirichlet energy goes to zero.
Definition 2.8.
A boundary end $\{E_{k}\}_{k=0}$ with length scale $L>0$ is nonparabolic if for every $k$, $E_{k}$ is an unbounded component of $M\backslash B_{kL+R_{0}}$ with nonparabolic boundary portion containing in $\partial M$ and $E_{k+1}\subset E_{k}$. Moreover, $E_{k}$ is nonparabolic.
An interior end is defined similarly.
3. One nonparabolic end for minimal hypersurface
In this section, following [Li-Wang-nonnegatively-curved-manifold], we prove that in a manifold $(X^{n+1},g)$ with $\operatorname{biRic}_{X}\geqslant 0$, the number of nonparabolic ends of a weakly stable minimal hypersurface $(M^{n},g)$ is at most one. This fact is known to experts. We include some details and adjustments for convenience of readers.
Theorem 3.1.
Let $(X^{n+1},g)$ be a complete manifold with bounded geometry and $\operatorname{biRic}_{X}\geqslant 0$. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete two-side weakly stable minimal immersion. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Then $M\backslash K$ admits at most one nonparabolic unbounded component. In particular, $M$ has at most one nonparabolic end.
We establish some preliminary results which are crucial in the proof of Theorem 3.1. In the sequel, we assume that the ambient manifold $(X^{n+1},g)$ always has bounded geometry.
3.1. Schoen–Yau inequality
The following inequality of Schoen and Yau is well known in nonnegative sectional curvature. Note that in dimension 4, Chodosh, Li and Stryker [chodosh-li-stryker] generalized it under assumption $\operatorname{Ric}^{X}_{2}\geqslant 0$. They also mentioned that in higher dimension $n+1$, $\operatorname{Ric}^{X}_{n-1}\geq 0$ plays the same role. The same proof works under the assumption $\operatorname{biRic}_{X}\geqslant 0$; see also [Cheng-Cheung-Zhou-CMC][Lemma 3.2] for the generalized inequality on CMC hypersurfaces. Note that $\operatorname{Ric}^{X}_{n-1}\geq 0$ implies $\operatorname{biRic}_{X}\geq 0.$ It would be interesting to understand more about these two assumptions.
Lemma 3.2.
Let $(X^{n+1},g)$ be a complete manifold with $\operatorname{biRic}_{X}\geqslant 0$. Let $M^{n}\hookrightarrow(X^{n+1},g)$ be a complete two-side weakly stable minimal immersion. Let $u$ be a harmonic function on $M$. Then
(3.1) $\frac{1}{n}\int_{M}\phi^{2}|A_{M}|^{2}|\nabla u|^{2}+\frac{1}{n-1}\int_{M}\phi^{2}|\nabla|\nabla u||^{2}\leqslant\int_{M}|\nabla\phi|^{2}|\nabla u|^{2}$
for any compactly supported, non-negative function $\phi\in W^{1,2}(M)$ satisfying
$\int_{M}\phi|\nabla u|=0.$
3.2. Infinite volume
The purpose of this subsection is to recall the following result which is first proved for complete orientable minimal hypersurface in $\mathbb{R}^{n+1}$ [Cao-Shen-Zhu-infinitevolume].
Theorem 3.3 ([Frensel-CMC]Theorem 1).
Let $M^{n}$ be a complete, noncompact manifold, and let $M^{n}\hookrightarrow X^{n+1}$ be an isometric immersion with bounded mean curvature. If $X^{n+1}$ has bounded geometry, the volume of $M^{n}$ is infinite.
3.3. One nonparabolic end
For contradiction, suppose that $\{E_{k}\}$ and $\{\widetilde{E}_{k}\}$ are distinct nonparabolic ends of $M$ adapted to a point $p\in M$ with length scale $L$. Since the ends are distinct, there is a $k_{0}\in\mathbb{N}$ such that $E_{k_{0}}\cap\widetilde{E}_{k_{0}}=\emptyset$.
To rule out two nonparabolic ends, we produce a non-constant harmonic function; see [chodosh-li-stryker][Lemma 4.3] for detailed proof.
Lemma 3.4.
Let $M^{n}$ be a complete Riemannian manifold. Let $K\subset M$ be a compact subset of $M$ with smooth boundary. Suppose that $M\backslash K$ has at least two nonparabolic unbounded components. Then there exists a non-constant harmonic function with finite Dirichlet energy on $M$.
We now turn to the proof of Theorem 3.1.
Proof of Theorem 3.1.
The proof follows from [Cheng-Cheung-Zhou-CMC][Theorem 3.1]. We sketch it here.
Suppose $E$ and $\widetilde{E}$ are nonparabolic unbounded components of $M\backslash K$. By Lemma 3.4, there exists a nonconstant harmonic function $u$ with finite Dirichlet energy. There exists a point $p\in M$ such that
$0<\int_{B_{1}(p)}|\nabla u|^{2}\leq\int_{B_{R}(p)}|\nabla u|^{2}\leqslant{C}\int_{\partial B_{R}(p)}|\nabla u|.$
Applying the co-area formula yields
$\int_{B_{R}(p)}|\nabla u|\geq C_{1}(R-1),$
which implies that $\int_{M}|\nabla u|=\infty$.
Let $\rho_{0}$ be a smoothing of $d_{M}(p,\cdot)$ satisfying $|\nabla\rho_{i}|<2$. We define
$\phi_{1}(R_{0},R):=\left\{\begin{array}[]{ll}1&x\in\overline{B}_{R_{0}}(p)\\ \frac{R_{0}+R-\rho_{0}(x)}{R}&x\in\overline{B}_{R_{0}+R}(p)\backslash B_{R_{0}}(p)\\ 0&x\in M\backslash B_{R_{0}+R}(p),\end{array}\right.$
and
$\phi_{2}(R_{0},R):=\left\{\begin{array}[]{ll}0&x\in\overline{B}_{R_{0}+R}(p)\\ \frac{R_{0}+R-\rho_{0}(x)}{R}&x\in\overline{B}_{R_{0}+2R}(p)\backslash B_{R_{0}+R}(p)\\ -1&x\in B_{R_{0}+2R+R_{1}}(p)\backslash B_{R_{0}+2R}(p)\\ \frac{\rho_{0}(x)-(R_{0}+3R+R_{1})}{R}&x\in\overline{B}_{R_{0}+3R+R_{1}}(p)\backslash B_{R_{0}+2R+R_{1}}(p)\\ 0&x\in M\backslash B_{R_{0}+3R+R_{1}}(p).\end{array}\right.$
Consider $\phi(t,R_{0},R)=\phi_{1}(R_{0},R)+t\phi_{2}(R_{0},R)$. Observe that choosing $R_{1}$ sufficiently large, we have
$\int_{M}\phi(0,R_{0},R)>0,\quad\int_{M}\phi(1,R_{0},R)<0.$
Therefore, there exists some $t_{0}\in(0,1)$ depending on $R_{0}$ and $R$ such that
$\int_{M}\phi(t_{0},R_{0},R)|\nabla u|=0.$
Applying Lemma 3.2, direct calculation yields
$\begin{split}&\frac{1}{n}\int_{B_{R_{0}}(p)}|A_{M}|^{2}|\nabla u|^{2}+\frac{1}{n-1}\int_{B_{R_{0}}(p)}|\nabla|\nabla u||^{2}\\ &\leqslant\int_{M}|\nabla\phi_{1}|^{2}|\nabla u|^{2}+t^{2}_{0}\int_{M}|\nabla\phi_{2}|^{2}|\nabla u|^{2}\\ &\leqslant\frac{4}{R^{2}}\int_{M}|\nabla u|^{2},\end{split}$
where we use $\langle\nabla\phi_{1},\nabla\phi_{2}\rangle=2$ in the first inequality. Letting $R\to\infty$, we conclude that
$|A_{M}|^{2}|\nabla u|^{2}\equiv 0,\quad|\nabla|\nabla u||^{2}\equiv 0.$
Note that $|\nabla u|$ is constant and $u$ is non-constant. We have that $|\nabla u|\neq 0$ everywhere. Theorem 3.3 implies that
$\int_{\Sigma}|\nabla u|^{2}=\operatorname{Vol}(M)|\nabla u|^{2}=\infty,$
which contradicts with the finiteness of Dirichlet energy. ∎
4. CMC hypersurfaces in $5$-manifolds with positive $k\operatorname{-triRic}$
Combining Lemma 2.3, Lemma 2.4 and Theorem 3.3, we can claim that $M^{4}$ is actually stable by [Barbosa-Berard-twisted-eigenvalue][Theorem 3.5]. Theorefore, without loss of generality, we can assume that $M^{4}$ is simply connected by looking at its universal cover. The parabolic ends of $M$ do not affect the construction of cut-off function on $M$ as we will see in the main proof (see also [chodosh-li-stryker][Remark 5.1]), thus we only need to deal with the nonparabolic end.
In this section, we adapt the argument from [chodosh-li-stryker] to prove a diameter bound that uses the theory of so-called balanced warped $\mu$-bubbles for stable minimal hypersurfaces; see also [Gromov19, chodosh-li-bubble] for relevant statements.
4.1. Balanced $\mu$-bubbles
We recall some preliminary results on existence and stability for balanced warped $\mu$-bubbles. Given a Riemannian $n$-manifold $(M^{n},g)$ with boundary, $n\leqslant 7$, we assume that a choice of labeling the components of $\partial M$ is $\partial M=\partial_{-}M\cup\partial_{+}M$ where neither of them are empty. For a fixed smooth function $u>0$ on $M$ and a smooth function on $\mathring{M}$ with $h\to\pm\infty$ on ±$\partial_{\pm}M$, consider the following balanced warped area functional, $k\in\mathbb{N}$.
[…]
for all Caccioppoli sets $\Omega$ with $\Omega\Delta\Omega_{0}\Subset\mathring{M}$ and positive $k$, where the reference set $\Omega_{0}$ with smooth boundary satisfies
$\partial\Omega_{0}\subset\mathring{M},\quad\partial_{+}M\subset\Omega_{0}.$
We will call $\Omega$ minimizing $\mathcal{A}_{k}$ in this class a balanced warped $\mu$-bubble, or just $\mu$-bubble for short.
Remark 4.1.
$\mathcal{A}_{k}(\Omega)$ is equivalent to the functional
[…]
under the conformal change $\widetilde{g}=u^{\frac{2k}{n-1}}g$. The critical point of $\widetilde{\mathcal{A}}$ is a hypersurface with the mean curvature $\tilde{H}=hu^{-\frac{k}{n-1}}$. This is equivalent to the following Lemma 4.3. We find that it is more convenient to work with $\mathcal{A}_{k}(\Omega)$ directly instead of $\widetilde{\mathcal{A}}(\Omega)$ under the conformal change.
4.1.1. Existence of balanced $\mu$-bubbles
The existence and regularity of a minimizer of $\mathcal{A}_{k}$ among all Caccioppoli sets (without any equivarient assumptions) was first claimed by Gromov in [Gromov19], and was rigorously carried out by Zhu in [Zhu-width-estimate]. We omit the detailed proof here and refer the readers to [chodosh-li-bubble].
Proposition 4.2.
There exists a smooth minimizer $\Omega$ for $\mathcal{A}_{k}$ such that $\Omega\Delta\Omega_{0}$ is compactly contained in the interior in $M$.
4.1.2. Stability
We list the first and second variation for a balanced warped $\mu$-bubble without proof; see [chodosh-li-bubble] for detailed calculation.
Lemma 4.3 (The first variation).
If $\Omega_{t}$ is a smooth $1$-parameter family of regions with $\Omega_{0}=\Omega$ and the normal speed at $t=0$ is $\psi$, then
[…]
where $\nu$ is the outwards pointing unit normal and $H$ is the scalar mean curvature of $\partial\Omega_{t}$. In particular, a $\mu$-bubble $\Omega$ satisfies
$H=-ku^{-1}\langle\nabla^{M}u,\nu\rangle+h,$
along $\partial\Omega$.
Lemma 4.4 (The second variation).
If $\Omega_{t}$ is a smooth $1$-parameter family of regions with $\Omega_{0}=\Omega$ and the normal speed at $t=0$ is $\psi$, then
$\begin{split}\frac{d^{2}}{dt^{2}}\bigg{|}_{t=0}\mathcal{A}_{k}(\Omega)&=\int_{\Sigma}|\nabla^{\Sigma}\psi|^{2}u^{k}-\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)\right)\psi^{2}u^{k}\\ &+\int_{M}k(k-1)u^{k-2}\left(\nabla_{\nu}u\right)^{2}\psi^{2}\\ &+\int_{\Sigma}ku^{k-1}\left(\Delta_{M}u-\Delta_{\Sigma}u\right)-\left(\nabla_{\nu}(u^{k}h)\right)\psi^{2}.\end{split}$
4.2. Diameter bound
We now turn to the diameter bound. We will need the following proposition.
Proposition 4.5.
Let $A=(a_{ij})$ be a $n\times n$ symmetric matrix and denote $H=tr(A)$, we have
$k|A|^{2}-a_{11}a_{nn}-\sum_{i=1}^{n}a_{1i}^{2}-\sum_{i=2}^{n}a_{in}^{2}+H(a_{11}+a_{nn})\geq\frac{2k^{2}+k-n+2}{2nk-3n+6}H^{2}$
for any $k\geq\frac{3}{2}.$ Equality holds if and only if
$a_{11}=a_{nn}=-\frac{2(n-2k-2)}{2nk-3n+6}H,\ \ a_{ii}=\frac{2k-1}{2nk-3n+6}H,\ \text{for}\ i\neq 1,n$
and $a_{ij}=0$ for any $i\neq j.$
Proof.
Since
$|A|^{2}\geq(a_{11}^{2}+a_{nn}^{2})+\sum_{i=2}^{n-1}a_{ii}^{2}+2(\sum_{i=2}^{n}a_{1i}^{2}+\sum_{i=2}^{n-1}a_{in}^{2})$
By Cauchy–Schwarz inequality, we have
$k|A|^{2}-a_{11}a_{nn}-\sum_{i=1}^{n}a_{1i}^{2}-\sum_{i=2}^{n}a_{in}^{2}+H(a_{11}+a_{nn})$
$\geq(k-1)(a_{11}^{2}+a_{nn}^{2})-a_{11}a_{nn}+\frac{k}{n-2}(H-a_{11}-a_{nn})^{2}$
$+(2k-1)(\sum_{i=2}^{n}a_{1i}^{2}+\sum_{i=2}^{n-1}a_{in}^{2})+H(a_{11}+a_{nn})$
$\geq(k-\frac{3}{2})(a_{11}^{2}+a_{nn}^{2})+\frac{k}{n-2}H^{2}-\frac{2kH}{n-2}(a_{11}+a_{nn})+\frac{k}{n-2}(a_{11}+a_{nn})^{2}$
$+H(a_{11}+a_{nn})$
$\geq(\frac{k}{2}-\frac{3}{4}+\frac{k}{n-2})(a_{11}+a_{nn})^{2}+(1-\frac{2k}{n-2})H(a_{11}+a_{nn})+\frac{k}{n-2}H^{2}.$
The conclusion follows from calculating the minimum. When the equality holds, $a_{11}=a_{nn}$ and $a_{ii}$ are all equal for $i=2,\cdots,n-1$. ∎
We first prove a diameter bound for the warped $\mu$-bubble.
Lemma 4.6.
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold. Let $M^{4}\hookrightarrow(X^{5},g)$ be a (strongly) stable two-sided CMC hypersurface with compact boundary $\partial M$ and of the mean curvature $H_{M}$. Suppose
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. There are universal constants $L>0$ and $c>0$ such that if there is a $p\in M$ with $d_{M}(p,\partial M)>\frac{L}{2}$, then in $M$ there is a relatively open set $\Omega\subset B^{M}_{L/2}(\partial M)$ and a smooth closed $3$-dimensional Riemannian manifold $\Sigma^{3}$ (possibly disconnected) such that $\partial\Omega=\Sigma\cup\partial M$ and each component of $\Sigma$ has diameter at most $c$.
Proof.
Let $\varphi_{0}$ be a smoothing of the function $d(\partial M,\cdot)$ so that $|\nabla\varphi_{0}|\leqslant 2$ and $\varphi_{0}\big{|}_{\partial M}\equiv 0$. Let $\epsilon\in(0,\frac{1}{2})$ so that $\epsilon$ and $\frac{4}{\sqrt{2\varepsilon_{0}}}\pi+2\epsilon$ are regular values of $\varphi_{0}$.
Define
$\varphi=\frac{\varphi_{0}-\epsilon}{\frac{4}{\sqrt{2\varepsilon_{0}}}+\frac{\epsilon}{\pi}}-\frac{\pi}{2}.$
$\Omega_{1}=\{x\in M:-\frac{\pi}{2}<\varphi<\frac{\pi}{2}\}$, and $\Omega_{0}=\{x\in\Sigma^{{}^{\prime}}:-\frac{\pi}{2}<\varphi\leqslant 0\}$. Clearly $|\operatorname{Lip}(\varphi)|\leqslant\frac{\sqrt{2\varepsilon_{0}}}{2}$, and we define
$h=-\sqrt{2\varepsilon_{0}}\tan(\varphi).$
We compute that
$\nabla h=-\sqrt{2\varepsilon_{0}}\left(1+\tan^{2}(\varphi)\right)\nabla\varphi=-\sqrt{2\varepsilon_{0}}\left(1+\frac{h^{2}}{2\varepsilon_{0}}\right)\nabla\varphi,$
so we have
$|\nabla h|\leqslant\varepsilon_{0}\left(1+\frac{h^{2}}{2\varepsilon_{0}}\right).$
We now consider the functional $\mathcal{A}_{k}$ on $M^{4}$, where the weight function $u$ satisfies
$\begin{split}-\Delta_{M}u&=\left(|A_{M}|^{2}+\operatorname{Ric}_{X}(\eta,\eta)\right)u~{}~{}\mbox{in}~{}~{}\mathring{M},\\ u&=1~{}~{}\mbox{on}~{}~{}\partial M;\end{split}$
see [Wu23, Hong24] for more details on the construction of $u$. We assume that $\Omega$ is the minimizer of $\mathcal{A}_{k}$ and $\Sigma$ is a connected component of $\partial\Omega\backslash\partial M$. $\nu$ is the unit normal vector to $\Sigma$.
To eliminate the term $\psi^{2}hu^{k-1}\nabla_{\nu}u$ in the second variation of $\Sigma$, note that by Lemma 4.3
$u^{k}H_{\Sigma}^{2}\psi^{2}=k^{2}u^{k-2}\left(\nabla_{\nu}u\right)^{2}\psi^{2}+u^{k}h^{2}\psi^{2}-2ku^{k-1}h\psi^{2}\nabla_{\nu}u.$
Since $k\leqslant 2$, plugging into the stability inequality yields
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\geqslant&\int_{\Sigma}\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-\frac{1}{2}H_{\Sigma}^{2}\right)u^{k}\psi^{2}\\ &\int_{\Sigma}-ku^{k-1}\psi^{2}\Delta_{M}u+\psi^{2}u^{k}\nabla_{\nu}h+\frac{1}{2}u^{k}h^{2}\psi^{2}.\end{split}$
We next deal with the curvature term
(4.1) $-ku^{-1}\Delta_{M}u+\left(|A_{\Sigma}|^{2}+\operatorname{Ric}^{M}(\nu,\nu)-\frac{1}{2}H_{\Sigma}^{2}\right).$
We select a local orthonormal frame $\{e_{i}\}_{i=1}^{3}$ on $\Sigma^{3}$ and extend it on $X^{5}$ by assigning $\nu=e_{4}$ and $\eta=e_{5}$. By Gauss equation, we have
$\begin{split}\operatorname{Ric}^{\Sigma}(e_{1},e_{1})&=\sum_{i=1}^{3}\operatorname{Rm}^{\Sigma}_{i11i}=\sum_{i=1}^{3}\left(\operatorname{Rm}^{M}_{i11i}+A^{\Sigma}_{ii}A^{\Sigma}_{11}-\left(A^{\Sigma}_{i1}\right)^{2}\right)\\ &=\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}\\ &=\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{M}_{i44i}-\operatorname{Ric}^{M}(\nu,\nu)+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}.\end{split}$
Therefore,
$\begin{split}k\operatorname{Ric}^{X}&(\eta,\eta)+\sum_{i=1}^{3}\operatorname{Rm}^{M}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{M}_{i44i}\\ =&k\sum_{i=1}^{5}\operatorname{Rm}^{X}_{i55i}+\sum_{i=1}^{3}\left(\operatorname{Rm}^{X}_{i11i}+A^{M}_{ii}A^{M}_{11}-\left(A^{M}_{i1}\right)^{2}\right)\\ &+\sum_{i=1}^{4}\left(\operatorname{Rm}^{X}_{i44i}+A^{M}_{ii}A^{M}_{44}-\left(A^{M}_{i4}\right)^{2}\right)\\ =&k\sum_{i=1}^{5}\operatorname{Rm}^{X}_{i55i}+\sum_{i=1}^{3}\operatorname{Rm}^{X}_{i11i}+\sum_{i=1}^{4}\operatorname{Rm}^{X}_{i44i}\\ &+(H_{M}-A^{M}_{44})A^{M}_{11}-\sum_{i=1}^{4}\left(A^{M}_{i1}\right)^{2}+(A^{M}_{41})^{2}+H_{M}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}\\ \geqslant&\lambda_{k}+(H_{M}-A^{M}_{44})A^{M}_{11}-\sum_{i=1}^{4}\left(A^{M}_{i1}\right)^{2}+(A^{M}_{41})^{2}+H_{M}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}.\end{split}$
Besides, by Proposition 4.5, all terms involving $A_{M}$ in (4.1) can be estimated as follows:
$k|A_{M}|^{2}-A^{M}_{11}A^{M}_{44}-\sum_{i=1}^{4}\left(A^{M}_{i4}\right)^{2}-\sum_{i=2}^{4}\left(A^{M}_{i4}\right)^{2}+H_{M}(A^{M}_{11}+A^{M}_{44})\geq\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}.$
It is not hard to check that
$|A_{\Sigma}|^{2}+H_{\Sigma}A^{\Sigma}_{11}-\sum_{i=1}^{3}\left(A^{\Sigma}_{i1}\right)^{2}\geqslant\frac{1}{2}H_{\Sigma}^{2}.$
In conclusion, we have the following inequality
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\geqslant&\int_{\Sigma}\left(\varepsilon_{0}-\operatorname{Ric}^{\Sigma}(e_{1},e_{1})\right)u^{k}\psi^{2}\\ &+\int_{\Sigma}\left(\varepsilon_{0}+\frac{1}{2}h^{2}+\nabla_{\nu}h\right)u^{k}\psi^{2}\end{split}$
We take $\psi=u^{-\frac{k}{2}}\phi$. Direct computation shows
$\nabla^{\Sigma}\psi=u^{-\frac{k}{2}}\nabla^{\Sigma}\phi-\frac{k}{2}u^{-\frac{k}{2}-1}\phi\nabla^{\Sigma}u,$
and
$u^{k}|\nabla^{\Sigma}\psi|^{2}=|\nabla^{\Sigma}\phi|^{2}+\frac{k^{2}}{4}|\nabla^{\Sigma}\log u|^{2}\phi^{2}-k\phi\langle\nabla^{\Sigma}\log u,\nabla^{\Sigma}\phi\rangle.$
Integration by parts and Cauchy–Schwarz inequality yield
$\begin{split}\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}&-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\\ &\leqslant\left(k\left(\frac{\varepsilon}{2}-1\right)+\frac{k^{2}}{4}\right)\int_{\Sigma}u^{-2}|\nabla^{\Sigma}u|^{2}\phi^{2}+\left(1+\frac{1}{2\varepsilon}\right)\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\\ &\leqslant\frac{4}{4-k}\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\end{split}$
by choosing $\varepsilon=2-\frac{k}{2}$. Choosing $e_{1}$ so that the minimum Ricci curvature scalar $\lambda_{\operatorname{Ric}}(\Sigma)=\operatorname{Ric}_{\Sigma}(e_{1},e_{1})$, we obtain that
$\begin{split}\frac{4}{4-k}\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}&\geqslant\int_{\Sigma}u^{k}|\nabla^{\Sigma}\psi|^{2}-ku^{k-1}\psi^{2}\Delta_{\Sigma}u\\ &\geqslant\int_{\Sigma}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)\phi^{2}.\end{split}$
Therefore, there is a smooth function $V\in C^{\infty}(\Sigma)$ so that
$V\geqslant\frac{4-k}{4}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)$
and
$\int_{\Sigma}|\nabla^{\Sigma}\phi|^{2}\geqslant\int_{\Sigma}V\phi^{2},$
for all $\phi\in C^{0,1}(\Sigma)$. Following the argument in [Chodosh-Li-Minter-Stryker][Theorem 5.1], the first eigenfunction of the operator $-\Delta_{\Sigma}-V$ is a smooth positive function $\theta$ that satisfies
$-\Delta_{\Sigma}\theta\geqslant\frac{4-k}{4}\left(-\lambda_{\operatorname{Ric}}(\Sigma)+\varepsilon_{0}\right)\theta.$
Let $\alpha=\frac{4}{4-k}$. Hence,
$\operatorname{Ric}^{(\theta,\alpha)}:=\operatorname{Ric}^{\Sigma}-\alpha^{-1}\left(\theta^{-1}\Delta_{\Sigma}\theta\right)g_{\Sigma}\geqslant\varepsilon_{0}g_{\Sigma}.$
Since $\alpha\leqslant 2$, we have by [shenyingyerugang][Corollary 1]
$\operatorname{diam}(\Sigma,g_{\Sigma})\leqslant\frac{2\pi}{\sqrt{2\varepsilon_{0}}}.$
Lemma 4.7.
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold. Let $M^{4}\hookrightarrow(X^{5},g)$ be a (strongly) stable two-sided CMC hypersurface of the mean curvature $H_{M}$ with nonsmooth boundary $\partial M=\partial_{0}M\cup\partial_{1}M$, and $\partial_{1}M$ intersects $\partial_{0}M$ at angle $(0,\pi/8)$. Suppose
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. There are universal constants $L>0$ and $c>0$ such that if there is a $p\in M$ with $d_{M}(p,\partial M)>\frac{L}{2}$, then in $M$ there is a relatively open set $\Omega\subset B^{M}_{L/2}(\partial_{1}M)$ and a smooth $3$-dimensional Riemannian manifold $\Sigma^{3}$ with boundary (possibly disconnected) such that $\Sigma=\partial\Omega\cap M^{\circ}$ and each component of $\Sigma$ has diameter at most $c$.
The proof of this Lemma follows the same procedure as Lemma 4.6 except the warped $\mu$-bubble has prescribed boundary. Refer the detailed proof of Theorem 4.4 in [Hong24] in ambient dimension $4$.
4.3. Almost linear volume growth
We show the almost linear growth.
Lemma 4.8 (Almost linear volume growth of a boundary or interior end).
Let $(X^{5},g)$ be a complete $5$-dimensional Riemannian manifold with bounded geometry and let $M^{4}\hookrightarrow(X^{5},g)$ be a simply-connected (strongly) stable CMC immersion (two-sided if it is minimal). Suppose $\operatorname{biRic}_{X}\geqslant 0$ and
$\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\varepsilon_{0}$
for a positive constant $\varepsilon_{0}$ and some $k\in[\frac{3}{2},2]$, where $\lambda_{k}$ is the infimum of the $k$-triRicci curvature. Let $\{E_{l}\}_{l\in\mathbb{N}}$ be an end of $M$ adapted to $p\in M$ with length scale $L$, where $L$ is the constant from Lemma 4.6 and Lemma 4.7. Let $M_{l}:=E_{l}\cap B_{(l+1)L}(p)$. Then there is a universal constant $C>0$ such that
$\operatorname{Vol}_{M}(M_{l})<C$
for all $l$ sufficiently large.
Proof.
The proof is very similar to that of [chodosh-li-stryker, Lemma 5.4]. For completeness, we sketch it here.
We apply Lemma 4.6 to the end $E_{l}$, which supplies an open set $\Omega_{l}$ in $B_{L/2}(\partial E_{l})\cap E_{l}$ with $\partial E_{l}\subset\partial\Omega_{l}$. Note that $\Omega_{l}\subset M_{l}$ and $\partial E_{l+1}\cap\overline{\Omega_{l}}=\emptyset$ by construction. Denote $\Sigma^{(i)}_{l}$ the components of $\Sigma_{l}:=\partial\Omega_{l}\backslash\partial E_{l}$.
As claimed in [chodosh-li-stryker, Lemma 5.4], $\overline{M_{l}}$ is connected and $\Sigma^{(i)}_{l}$ separates $\partial E_{l}$ and $\partial E_{l+1}$. By Lemma 4.6, we have $\operatorname{diam}(\Sigma_{l})\leqslant c$.
We claim that there is a constant $D>0$ independent of $l$ such that $\operatorname{diam}(M_{l})<D$. Let $q$ and $r$ be any points in $M_{l}$. The radial geodesic connecting them intersects $\partial E_{l-1}$ and $\partial E_{l}$. By the choice of the components of $\{\Sigma_{l}\}$, an arc of this curve of length at most $2L$ joins $q$ to $\Sigma_{l-1}$. The same argument applies to $r$. Therefore, the diameter of $M_{l}$ is bounded by $D:=4L+c$.
By Lemma 2.4, we know that $M$ has bounded second fundamental form (e.g. consider the comapct immersion of $B_{M}(q,1)$ for any $q\in M$). By Lemma 2.3, it implies that $\operatorname{Ric}_{M}$ is bounded from below. Hence, by Bishop–Gromov volume comparsion, there is a constant $C>0$ so that
$\operatorname{Vol}(B_{M}(x,D))\leqslant C$
for all $x\in M$. Since $M_{l}\subset\operatorname{Vol}(B_{M}(x,D))$ for any $x\in M_{l}$, the lemma follows. ∎
5. Proof of main results
5.1. Proof of Theorem 1.9
The conclusion follows from the argument in [chodosh-li-stryker] by constructing compactly supported test function. While, for completeness and independent interest (to prove Proposition 1.13), we provide the detailed construction of balanced test functions for weak stability. To do it, we need to decompose $M$ into convenient building-blocks (see [chodosh-li-stryker]).
Let $M^{4}$ be a complete, simply connected, Riemannian $4$-manifold such that the conclusions of Theorem 3.1 hold. Let $\{E_{l}\}$ be the nonparabolic end of $M$ adapted to $p\in M$ with length scale $L$.
We now decompose $M$ to deal with the nonparabolic end. Let $l_{0}\geqslant 1$. Then
  • $M_{l}:=E_{l}\cap B_{(l+1)L}(p)$ for all $l$.
  • $\{P^{(\alpha)}_{l_{0}}\}_{\alpha=1}^{n_{l_{0}}}$ are the components of $M\backslash\overline{B}_{l_{0}L}(p)$ besides $E_{l_{0}}$.
  • $\{P^{(\alpha)}_{l}\}_{\alpha=1}^{n_{l}}$ are the components of $E_{l-1}\backslash\overline{B}_{lL}(p)$ besides $E_{l}$ for $l>l_{0}$.
The third part aims to pack the parabolic ends attached to the nonparabolic end $E_{l}$.
Figure 1. Decomposition of $M$
Note that for fixed $i\geq 1$, we have the decomposition
$M=\overline{B}_{l_{0}L}(p)\cup\bigcup_{l=l_{0}}^{l_{0}+i-1}\left(\overline{M}_{l}\cup\bigcup_{\alpha=1}^{n_{l}}\overline{P}_{l}^{(\alpha)}\right)\cup\left(\overline{E}_{l_{0}+i-1}\backslash B_{(l_{0}+i)L}(p)\right).$
By Theorem 3.1, $P_{l}^{(\alpha)}$ is either bounded or $P_{l}^{(\alpha)}\backslash K$ is parabolic for all compact $K\subset M$ such that $P_{l}^{(\alpha)}\backslash K$ has smooth boundary. It is important to keep in mind that as $l\rightarrow\infty$, $n_{l}$ may not be bounded.
To prove the main theorem, we recall the following fact.
Proposition 5.1.
For any $l$ and $\alpha\in[1,n_{l}]$ and for any small $\epsilon>0$, there exists a positive compactly supported Lipschitz function $u_{l,\alpha}$ on $P_{l}^{(\alpha)}$ such that
$u_{l,\alpha}|_{\partial P_{l}^{(\alpha)}}=1\ \text{and}\ \ \int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}<\epsilon.$
This follows from the definition of the parabolicity of ends. See [peterwangspectrum] or [chodosh-li-stryker, Proposition 3.4] for the detailed explanation.
Proof of Theorem 1.9.
Suppose that a two-sided weakly stable minimal hypersurface $M^{4}$ exists in $(X^{5},g)$ of Theorem 1.9, let $\{E_{l}\}$ be the nonparabolic end of $M$. We use the deocomposition of $M$ constructed above.
For each $l$, let $\rho_{l}$ be a smooth function on $\overline{M}_{l}$ such that $|\nabla\rho_{l}|\leqslant 2$,
$\rho_{l}\big{|}_{\partial E_{l}}\equiv lL,\quad\rho_{l}\big{|}_{\partial M_{l}\backslash\partial E_{l}}\equiv(l+1)L.$
For instance, we can take $\rho_{l}$ to be a smoothing of the distance function from $p$.
Since $M$ is weakly stable, outside a compact subset $\Omega$, $M$ is strongly stable. We can choose large $l_{0}$ such that $B_{l_{0}L}(p)$ contains $\Omega$. To prove the theorem, on the contrary we assume that $\omega(x)=|A_{M}|^{2}(x)+\operatorname{Ric}_{X}(\eta,\eta)$ is positive at some point $q$, then $\omega$ is positive in a neighborhood $\Omega_{q}$ of $q$ by continuity. We also assume that $\Omega_{q}$ is contained in $B_{l_{0}L}(p)$, this is achievable since $p$ and ł$\l_{0}$ are free to choose.
Let
$\int_{B_{l_{0}L}(p)}w\geq T$
where $T>0$ is fixed positive value.
Let $\phi_{1}:[l_{0}L,(l_{0}+2i)L]\to[-1,1]$ be the linear function defined as
$\phi_{1}(t)=1+\frac{1}{i}(l_{0}-\frac{t}{L}).$
Define a compactly supported Lipschitz function on $M$ as follows
$\varphi_{1}(x)=\begin{cases}1\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p),\\ \phi_{1}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}\leq l<l_{0}+i,\\ \phi_{1}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}\leq l<l_{0}+i,\\ 0&\ \ \text{otherwise}.\end{cases}$
Define a Lipschitz function on $M$ as follows
$\varphi_{2}(x)=\begin{cases}0,\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p)\cup\left(\cup_{l=l_{0}}^{l_{0}+i-1}\overline{M_{l}}\right)\cup\left(\cup_{l=l_{0}}^{l_{0}+i-1}\overline{P}_{l}^{(\alpha)}\right),\\ \phi_{1}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}+i\leq l<l_{0}+2i,\\ \phi_{1}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+i\leq l<l_{0}+2i,\\ -u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+2i\leq l<l_{0}+3i+a,\\ -1&\text{otherwise}.\end{cases}$
Note that the two functions $\varphi_{1},\varphi_{2}$ have different supports on $M$.
Let $\phi_{2}:[(l_{0}+2i+a)L,(l_{0}+3i+a)L]\to[0,1]$ be the linear function defined as
$\phi_{2}(t)=-2+\frac{1}{i}(\frac{t}{L}-l_{0}-a).$
Define a Lipschitz function on $M$ as follows
$\varphi_{a}(x)=\begin{cases}0\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p)\cup\left(\cup_{l=l_{0}}^{l_{0}+2i+a-1}\overline{M_{l}}\right)\cup\left(\cup_{l=l_{0}}^{l_{0}+2i+a-1}\overline{P}_{l}^{(\alpha)}\right),\\ \phi_{2}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}+2i+a\leq l<l_{0}+3i+a,\\ \phi_{2}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+2i+a\leq l<l_{0}+3i+a,\\ 1&\ \ \text{otherwise}.\end{cases}$
Thus we can sum $\varphi_{2}$ and $\varphi_{a}$ to get a compactly supported Lispchtiz function. Let $\varphi_{s,a}=\varphi_{1}+s(\varphi_{2}+\varphi_{a})$ and we can easily see that $\varphi_{t}$ has compact support in $M$. We calculate
$\int_{M}\varphi_{1,a}$ $=\int_{M}\varphi_{1}+\int_{M}\left(\varphi_{2}+\varphi_{a}\right)$
$\leq\int_{M}\varphi_{1}-\sum_{l=l_{0}+2i}^{l_{0}+2i+a-1}\sum_{\alpha=1}^{n_{l}}\int_{P_{l}^{\alpha}}u_{l,\alpha}-\sum_{l=l_{0}+2i}^{l_{0}+2i+a-1}|M_{l}|.$
The first term of right hand side is fixed value. Thus we are able to choose large $a$ such that $\int_{M}\varphi_{1,a}<0$, due to the fact that $M$ has infinite volume. Denote such value by $a_{0}.$
Moreover, we know that $\int_{M}\varphi_{0,a_{0}}>0$. Since $\int_{M}\varphi_{s,a_{0}}$ is continuous in $s$, there exists $s_{0}\in(0,1)$ such that $\int_{M}\varphi_{s_{0}.a_{0}}=0$. Hence, it follows from the part regarding the interior ends of Lemma 4.8 and the direct computation that
$\int_{M}$ $|\nabla\varphi_{s_{0},a_{0}}|^{2}\leq\int_{M}|\nabla\varphi_{1}|^{2}+\int_{M}|\nabla\varphi_{2}|^{2}+\int_{M}|\nabla\varphi_{a_{0}}|^{2}+2\int_{M}\langle\nabla\varphi_{2},\nabla\varphi_{a_{0}}\rangle$
$=\sum_{l=l_{0}}^{l_{0}+2i-1}\int_{M_{l}}(\phi_{1}^{{}^{\prime}})^{2}|\nabla\rho_{l}|^{2}+\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\int_{M_{l}}(\phi_{2}^{{}^{\prime}})^{2}|\nabla\rho_{l}|^{2}$
$+\sum_{l=l_{0}}^{l_{0}+2i-1}\sum_{\alpha=1}^{n_{l}}\phi_{1}^{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}-2\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\phi_{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}$
$+\sum_{l=l_{0}+2i}^{l_{0}+2i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}+\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\phi_{2}^{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}$
By Proposition 5.1, we set $\epsilon=\frac{1}{n_{l}(3i+a_{0})^{2}}$ in each sum. Therefore, we obtain
$\int_{M}|\nabla\varphi_{s_{0},a_{0}}|^{2}\leq 4C\frac{1}{(iL)^{2}}2i+4C\frac{1}{(iL)^{2}}i+\frac{1}{3i+a_{0}}\leq\frac{C}{i}.$
By weak stability, we obtain
00 0 $\leq Q(\varphi_{s_{0},a_{0}},\varphi_{s_{0},a_{0}})$
$=\int_{M}|\nabla\varphi_{s_{0},a_{0}}|^{2}-\int_{M}w\varphi_{s_{0},a_{0}}^{2}$
$\leq C/i-\int_{B_{l_{0}L}(p)}w$
$\leq C/i-T.$
Letting $i$ tend to the infinity leads to a contradiction. Thus $w(x)$ has to vanish identically. That is, $A_{M}\equiv 0$ and $\operatorname{Ric}_{X}(\eta,\eta)=0$ along $M$, completing the proof.
On the one hand, the above proof also works for weakly stable CMC hypersurface. On the other hand, as mentioned before it can be adjusted to prove Proposition 1.13.
5.2. Proof of Theorem 1.12
In this subsection, we prove Theorem 1.12. We emphasize that the mean curvature used in the sequel is the trace of the second fundamental form. First, we prove the following.
Theorem 5.2.
A complete constant mean curvature hypersurface $M^{4}$ in $(X^{5},g)$ with the bounded geometry satisfying universal subexponential end growth condition,
(5.1) $\frac{H_{M}^{2}}{4}+\inf_{X}\operatorname{Ric}_{X}\geq\delta,$
and
(5.2) $\inf_{X}k\operatorname{-triRic}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\epsilon_{0}$
for positive constants $\delta$, $\epsilon_{0}$ and $k\in[\frac{3}{2},2]$ as well as
$\operatorname{Ind}<\infty$
must be compact.
Proof.
The proof follows the same strategy in [Hong24] by the first author. For completeness we include details. Indeed, on the contrary, we assume that there exists a complete noncompact hypersurface $M$ in $X^{5}$ satisfying finite index condition, universal subexponential end growth condition and the curvature conditions in the theorem.
Since $M$ has finite index, in any end $E$ of $M$ (outside a large compact subset) there holds
[…]
$\geq\left(\frac{H_{M}^{2}}{4}+\inf_{X}\operatorname{Ric}_{X}\right)\int_{E}\varphi^{2}$
$\geq\delta\int_{E}\varphi^{2}$
for any $\varphi\in C^{\infty}_{c}(E)$. It follows from [peterwangspectrum] that if $E$ is a parabolic end, it satisfies exponential volume decay, thus has finite volume. However, in a Riemannian manifold with bounded geometry, any complete noncompact CMC hypersurface (in fact, any end of this hypersurface) has infinite volume. This contradiction implies that $M$ has no parabolic end. Thus all ends of $M$ are nonparabolic. Notice that the number of nonparabolic ends may not be finite.
Again for a finite index CMC hypersurface, there exists a compact subset such that the hypersurface is strongly stable outside (see [Fischer-Colbrie-On-complete-minimal]). Take a nonparabolic end $E$ with respect to the compact subset such that $E$ is stable. Consider its universal covering $\tilde{E}$, which is still a stable CMC hypersurface with boundary in $X^{4}$ satisfying subexponential end growth condition by the assumption.
On the one hand, by Lemma 5.1 in [Hong24], $\tilde{E}$ is nonparabolic. According to Corollary 3.6 and Lemma 3.7 in [Hong24], $\tilde{E}$ satisfies exponential volume growth:
$|\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})|\geq C\cdot e^{Cl}$
for a constant $C=C(\delta)$ depending on $\delta.$ On the other hand, by Lemma 4.8 each component of $\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})$ has uniform volume upper bound. It follows from the subexponential end growth condition that the volume of $\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})$ satisfies
$|\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})|=o(e^{cl})$
for any $c>0.$ This contradicts with the exponential volume growth. ∎
Proof of Theorem 1.12.
We only need to check the curvature conditions in Theorem 5.2. The first one leads to that $H_{M}>4$. The second one turns to finding the minimum of $H_{M}$ satisfying the followng inequality in terms of $k\in[\frac{3}{2},2]$:
$H_{M}^{2}>\frac{(8k-6)(4k+5)}{2k^{2}+k-2}.$
When $k=2,$ the right hand side achieves the minimum $65/4$
Appendix A Example
We provide a closed $5$-dimensional Riemannian manifold $(X^{5},g)$ with bounded geometry, satisfying
$\operatorname{Ric}_{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ 2\operatorname{-triRic}_{X}\geq\delta>0,$
while $\operatorname{Ric}_{3}<0$ somewhere.
Consider the Berger sphere: $\mathbb{S}^{5}=\overline{G}/\overline{H}$, where $\overline{G}=SU(3)\times\mathbb{S}^{1}$ and $\overline{H}=SU(2)\times\mathbb{S}^{1}$. We follow the notation in [Ziller-Bergersphere]. For the bi-invariant metric $\langle X,Y\rangle=-\frac{1}{2}\operatorname{tr}(XY)$, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is an $\operatorname{Ad}(H)$ invariant splitting and $\mathfrak{p}=\mathfrak{h}^{\perp}$ has an orthonormal basis: for $r\in\{1,2\}$,
$A=\frac{i}{\sqrt{3}}\left(E_{1,1}+E_{2,2}-2E_{3,3}\right),\quad e_{r}=E_{r,3}-E_{3,r},\quad f_{r}=i\left(E_{r,3}+E_{3,r}\right).$
Moreover, $\mathfrak{p}=\mathfrak{p}_{1}\oplus\mathfrak{p}_{2}$, where
$\mathfrak{p}_{1}:=\operatorname{span}\{e_{r},f_{r}\},\quad\mathfrak{p}_{2}=\mathbb{R}\cdot A.$
Let $D$ be a basis of the Lie algebra of $\mathbb{S}^{1}$. By [Ziller-Bergersphere][Theorem 3], the metric $g_{s}$ defined by
$g_{s}:=g|_{\mathfrak{p}_{1}\times\mathfrak{p}_{1}}+s^{2}\cdot g|_{\mathfrak{p}_{2}\times\mathfrak{p}_{2}}$
is naturally reductive with respect to the decomposition $\overline{\mathfrak{g}}=\mathfrak{h}\oplus\overline{\mathfrak{p}}_{s}$, where
$\overline{\mathfrak{p}}_{s}:=\operatorname{span}\{s^{2}A+(s^{2}-1)D,e_{r},f_{r}\}.$
We will abbreviate
$d_{s}=\frac{s^{2}A+(s^{2}-1)D}{s}.$
Thus $\{d_{s},e_{r},f_{r}\}$, $r\in\{1,2\}$, is a $g_{s}$-orthonormal basis of $\overline{\mathfrak{p}}_{s}$. Since $\operatorname{Ad}(\overline{H})$ maps any unit vector into a vector
$v_{\alpha}:=\cos\alpha d_{s}+\sin\alpha e_{1},$
without loss of generality, we can restrict ourselves to $v=v_{\alpha}$. For convenience, we set $\overline{e}_{1}=\sin\alpha d_{s}-\cos\alpha e_{1}$.
Following the calculation in [Ziller-Bergersphere], we list sectional curvatures of $(\mathbb{S}^{5},g_{s})$:
$\begin{split}\sec_{g_{s}}(v,\overline{e}_{1})=\frac{3}{4}s^{2},\quad&\sec_{g_{s}}(v,e_{2})=\sin^{2}\alpha+\frac{3s^{2}}{4}\cos^{2}\alpha,\\ \sec_{g_{s}}(v,f_{1})=\frac{3}{4}s^{2}+(4-3s^{2})\sin^{2}\alpha,\quad&\sec_{g_{s}}(v,f_{2})=\sin^{2}\alpha+\frac{3s^{2}}{4}\cos^{2}\alpha,\\ \sec_{g_{s}}(\overline{e}_{1},e_{2})=\cos^{2}\alpha+\frac{3s^{2}}{4}\sin^{2}\alpha,\quad&\sec_{g_{s}}(\overline{e}_{1},f_{1})=\frac{3}{4}s^{2}+(4-3s^{2})\cos^{2}\alpha,\\ \sec_{g_{s}}(\overline{e}_{1},f_{2})=\cos^{2}\alpha+\frac{3s^{2}}{4}\sin^{2}\alpha,\quad&\sec_{g_{s}}(e_{2},f_{1})=1,\\ \sec_{g_{s}}(e_{2},f_{2})=4-\frac{9}{4}s^{2},\quad&\sec_{g_{s}}(f_{1},f_{2})=1.\end{split}$
Elementary calculation yields
$\begin{split}\operatorname{Ric}_{g_{s}}(v,v)&=3s^{2}+\left(6-\frac{9}{2}s^{2}\right)\sin^{2}\alpha,\\ \operatorname{Ric}_{g_{s}}(\overline{e}_{1},\overline{e}_{1})&=6-\frac{3}{2}s^{2}+\left(\frac{9}{2}s^{2}-6\right)\sin^{2}\alpha\\ \operatorname{Ric}_{g_{s}}(e_{2},e_{2})=\operatorname{Ric}_{g_{s}}(f_{1},f_{1})&=\operatorname{Ric}_{g_{s}}(f_{2},f_{2})=6-\frac{3}{2}s^{2},\\ \operatorname{Ric}_{g_{s}}&\equiv 0,~{}~{}\mbox{otherwise.}~{}~{}\end{split}$
When $s^{2}\in(16/9,4]$, we have
$\inf\sec_{g_{s}}=4-\frac{9}{4}s^{2}<0,\quad\inf\operatorname{Ric}_{g_{s}}=6-\frac{3}{2}s^{2}\geq 0.$
Note that in general,
$\operatorname{Ric}_{g_{s}}(v,v)\leq\operatorname{Ric}_{g_{s}}(e_{2},e_{2})=\operatorname{Ric}_{g_{s}}(f_{1},f_{1})=\operatorname{Ric}_{g_{s}}(f_{2},f_{2})\leq\operatorname{Ric}_{g_{s}}(\overline{e}_{1},\overline{e}_{1}).$
Next, we list all the bi-Ricci curvatures:
$\begin{split}\operatorname{biRic}_{g_{s}}(v,\overline{e}_{1})&=6+\frac{3}{4}s^{2},\quad\operatorname{biRic}_{g_{s}}(v,f_{1})=6+\frac{3}{4}s^{2}+\left(2-\frac{3}{2}s^{2}\right)\sin^{2}\alpha,\\ \operatorname{biRic}_{g_{s}}(v,e_{2})&=\operatorname{biRic}_{g_{s}}(v,f_{2})=6+\frac{3}{4}s^{2}+\left(5-\frac{15}{4}s^{2}\right)\sin^{2}\alpha,\\ \operatorname{biRic}_{g_{s}}(\overline{e}_{1},e_{2})&=\operatorname{biRic}_{g_{s}}(\overline{e}_{1},f_{2})=11-3s^{2}+\left(-5+\frac{15}{4}s^{2}\right)\sin^{2}\alpha,\\ \operatorname{biRic}_{g_{s}}(\overline{e}_{1},f_{1})&=8-\frac{3}{4}s^{2}+\left(-2+\frac{3}{2}s^{2}\right)\sin^{2}\alpha,\quad\operatorname{biRic}_{g_{s}}(e_{2},f_{1})=11-3s^{2},\\ \operatorname{biRic}_{g_{s}}(e_{2},f_{2})&=8-\frac{3}{4}s^{2},\quad\operatorname{biRic}_{g_{s}}(f_{1},f_{2})=11-3s^{2}.\end{split}$
Therefore, when $s^{2}\in(16/9,11/3]$, we have $\operatorname{biRic}\geq 0$.
Recall that
$2\operatorname{-triRic}(x;y,z)=2\operatorname{Ric}(x,x)+\operatorname{biRic}(y,z)-\sec(x,y)-\sec(x,z).$
Therefore, if $02\operatorname{Ric}(x,x)-\sec(x,y)-\sec(x,z)>$ 0, we have $02\operatorname{-triRic}(x;y,z)>$ 0 automatically.
Direct calculation yields when $s^{2}\in(16/9,11/3]$
$\begin{split}2\operatorname{-triRic}(v;\overline{e}_{1},e_{2})=2\operatorname{-triRic}(v;\overline{e}_{1},f_{2})&\geq 11+\frac{3}{2}s^{2}+6\left(1-\frac{3}{4}s^{2}\right)>0,\\ 2\operatorname{-triRic}(v;\overline{e}_{1},f_{1})>0,\quad&2\operatorname{-triRic}(v;e_{2},f_{1})>0,\\ 2\operatorname{-triRic}(v;e_{2},f_{2})\geq 18-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(v;f_{1},f_{2})>0.\end{split}$
$\begin{split}2\operatorname{-triRic}(\overline{e}_{1};v,e_{2})=2\operatorname{-triRic}(\overline{e}_{1};v,f_{2})&\geq 17-3s^{2}>0,\\ 2\operatorname{-triRic}(\overline{e}_{1};v,f_{1})>0,\quad&2\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{1})>0,\\ 2\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{2})\geq 20-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(\overline{e}_{1};f_{1},f_{2})>0,\end{split}$
$\begin{split}2\operatorname{-triRic}(e_{2};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(e_{2};v,f_{1})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(e_{2};v,f_{2})>0,\quad&2\operatorname{-triRic}(e_{2};\overline{e}_{1},f_{1})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(e_{2};\overline{e}_{1},f_{2})>0,\quad&2\operatorname{-triRic}(e_{2};f_{1},f_{2})>0.\end{split}$
$\begin{split}2\operatorname{-triRic}(f_{1};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(f_{1};v,e_{2})>0,\\ 2\operatorname{-triRic}(f_{1};v,f_{2})\geq 21-6s^{2},\quad&2\operatorname{-triRic}(f_{1};\overline{e}_{1},e_{2})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(f_{1};\overline{e}_{1},f_{2})>0,\quad&2\operatorname{-triRic}(f_{1};e_{2},f_{2})\geq 20-\frac{15}{4}s^{2}>0.\end{split}$
To obtain $02\operatorname{-triRic}(f_{1};v,f_{2})>$ 0, we require $s^{2}<7/2<11/3$.
$\begin{split}2\operatorname{-triRic}(f_{2};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};v,e_{2})>0,\\ 2\operatorname{-triRic}(f_{2};v,f_{1})\geq 18-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};\overline{e}_{1},e_{2})>0,\\ 2\operatorname{-triRic}(f_{2};\overline{e}_{1},f_{1})\geq 20-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};e_{2},f_{1})>0.\end{split}$
In conclusion, when $216/9<s^{2}<7/$ 2, $(\mathbb{S}^{5},g_{s})$ has positive $\operatorname{Ric}$, $\operatorname{biRic}$ and $2\operatorname{-triRic}$ curvatures with some negative sectional curvature.
However, when $210/3<s^{2}<7/$ 2,
$\sec(\overline{e}_{1},f_{1})+\sec(v,f_{1})+\sec(e_{2},f_{1})=-\frac{3}{2}s^{2}+5<0,$
which implies that $\operatorname{Ric}_{3}<0$ somewhere. By the main theorem, such ambient manifold does not admit any two-sided complete noncompact minimal hypersurfaces.
Moreover, note that when $s^{2}\in(\frac{11}{3},4)$, $\operatorname{Ric}_{g_{s}}\geq 3\left(2-\frac{s^{2}}{2}\right)>0$, while $\operatorname{biRic}_{g_{s}}(e_{1},e_{2})=11-3s^{2}<0$ and $\inf\sec_{g_{s}}=4-\frac{9}{4}s^{2}<0$. It is interesting to understand whether there exists a two-sided complete noncompact stable minimal hypersurface in such manifold $(\mathbb{S}^{5},g_{s})$.